# Identification of fixations and saccades in eye-tracking data using adaptive threshold-based methods

### v2

@@ 4 Discussion and conclusion
This study compared three threshold-based classifiers, I-VT, I-AVT, and I-DT, on free viewing (Random Pixel) and visual search (Waldo) eye-tracking data, using EyeLink labels as a reference. Across both tasks, I-VT provided the highest overall agreement with the benchmark and the most favorable balance between fixation and saccade classification. Task structure influenced the optimal thresholds obtained by $K$-ratio minimization. The Waldo dataset required higher velocity and angular-velocity thresholds than the Random Pixel dataset, which reflects more variable and goal-directed oculomotor dynamics in visual search. At the same time, both datasets were dominated by fixations, with saccades forming only a small fraction of samples. This class imbalance means that accuracy alone is not a sufficient measure of performance. The class-wise metrics in the Results show that a classifier can achieve relatively high accuracy while failing to identify many saccades. This supports the use of F1-scores, confusion matrices, and explicit event proportions, in line with earlier discussions of evaluation practice in eye movement classification [zemblys].
@@
The $K$-ratio criterion provided a simple, task-specific method to tune thresholds without explicit modelling of the underlying velocity distributions. By minimizing the discrepancy between empirical and independent transition probabilities, it favored segmentation in which fixations and saccades formed temporally coherent states. This approach is complementary to more elaborate state-space models, such as hidden Markov models, which require stronger assumptions about observation statistics [10.1145/355017.355028]. Here, the $K$-ratio produced consistent thresholds across tasks and offered a transparent way to adapt classical I-VT, I-AVT, and I-DT schemes to raw eye-tracking data.
@@
The noise analysis confirmed that all three algorithms are sensitive to measurement noise. Even modest Gaussian perturbations led to a marked reduction in agreement with the EyeLink labels and to substantial shifts in event composition. In all cases, re-optimizing thresholds by $K$-ratio mitigated, but did not entirely prevent, the loss of performance. These findings agree with previous work emphasizing the importance of careful preprocessing and denoising in eye-tracking analysis [niehorster2020noise, niehorster2017denoising, LARSSON2015145]. From an applications perspective, the results suggest that I-VT remains a reasonable default when high-quality data are available, and the goal is to obtain a reliable first-order segmentation of fixations and saccades. I-AVT, despite its directional component, offered limited benefit in this context because of its strong under-detection of saccades. I-DT may still be useful in settings where long, stable fixations are of primary interest, such as some reading and static-scene paradigms [68d3bfb814cc4a61aae0131dd135c9ec], but it is less suitable for dynamic visual search. In all cases, practitioners should be aware that fixed thresholds calibrated on one task or recording setup may not transfer to others, and that auxiliary events such as microsaccades, smooth pursuits, and glissades remain largely unmodeled [martinezs].
@@
There are several natural directions for future work. One is to combine $K$-ratio tuning with more advanced denoising and interpolation schemes, which could extend the useful operating range of classical algorithms in noisy or mobile settings. Another is to embed $K$-ratio or related transition statistics into learning-based classifiers, for example, as regularization terms or diagnostic tools for model calibration [FRANCESCHIELLO2022106929, zemblys]. A third is to pursue more systematic study of inter-individual differences, with thresholds optimized at the individual level, to clarify how much of the observed variability is due to genuine oculomotor differences and how much is an artifact of algorithm choice [HESSELS2019100710].
@@
Declarations
  • Conflict of interest/Competing interests: The authors have no competing interests to declare that are relevant to the content of this article.
  • Data availability: Data sets used during the current study are available from the corresponding author on reasonable request.
@@ Acknowledgements
Supplementary results
@@
In this supplementary section, we provide detailed results from the noise robustness analysis. Table S2 reports the task-specific thresholds obtained by $K$-ratio minimization under noise-free conditions. These values served as the fixed (non-adaptive) baseline thresholds in subsequent experiments.
@@
Table S2: $K$-ratio-optimized thresholds for the three algorithms (zero added noise).
@@
Table S3: Classification degradation with fixed thresholds under increasing noise for I-VT algorithm. Note: Agr. = agreement with noise level 0 ($\sigma=0$) (%); Fix/Sac = sample counts; $\Delta$% = change in fixation proportion from baseline.
@@
Table S4: I-VT: Original vs. Adaptive threshold performance across noise levels. Note: Acc = Accuracy; F1F = F1-score (Fixation); F1S = F1-score (Saccade).
@@
Table S5: I-AVT: Original vs. Adaptive threshold performance across noise levels. See Tab. S4.
@@
Table S6: I-DT: Original vs. Adaptive threshold performance across noise levels. See Tab. S4.
@@
Figure S5: F1 scores for fixation and saccade classification across noise levels. Top row: F1-score for fixation detection. Bottom row: F1-score for saccade detection. Solid lines indicate original thresholds; dashed lines indicate adaptive thresholds.
Figure S6: Precision and recall for fixation detection across noise levels. Top row: Precision. Bottom row: Recall. Adaptive thresholds maintain higher precision at the cost of reduced recall at extreme noise levels.
Table S3 shows how classification performance degrades when optimal noise-free thresholds are held fixed under increasing Gaussian spatial noise ($\sigma=0$5050 px) for the I-VT algorithm. Agreement with the noise-free classification drops below 90% at $\sigma=2$ px for Random Pixel, while Waldo maintains agreement above 90% until $\sigma=5$ px. At high noise levels ($\sigma\geq 30$ px), both tasks approach chance-level agreement ($\sim$5–7%), with nearly all samples misclassified as saccades.
Tables S4-S6 compare accuracy and F1 scores between fixed thresholds (optimized at $\sigma=0$) and adaptive thresholds (re-optimized at each noise level) for all three algorithms across both tasks. I-DT with adaptive thresholds demonstrates the greatest robustness, maintaining accuracy above 0.81 even at $\sigma=50$ px, whereas I-AVT shows the steepest degradation, falling to 0.31 accuracy at extreme noise levels.
Figure S5 visualizes F1 scores for fixation and saccade detection across noise levels. Adaptive thresholds (dashed lines) preserve fixation F1 scores substantially better than fixed thresholds (solid lines), with I-DT maintaining fixation F1 >> 0.90 across all noise levels. However, saccade F1 scores decline for all algorithms under both threshold conditions, with adaptive I-DT showing particularly low saccade F1 (<< 0.11) at high noise. Figure S6 reveals the precision-recall trade-off underlying adaptive threshold optimization. Precision for fixation detection remains high across noise levels under adaptive thresholding, but recall decreases substantially at $\sigma>10$ px. This pattern indicates that adaptive thresholds achieve overall accuracy by conservatively classifying uncertain samples as fixations.

### v3

@@ 4 Discussion and conclusion
This study compared three threshold-based classifiers, I-VT, I-AVT, and I-DT, on free viewing (Random Pixel) and visual search (Waldo) eye-tracking data, using EyeLink labels as a reference. Because the EyeLink parser is itself a threshold-based classifier, our reported accuracies quantify agreement with a widely used commercial event-detection pipeline, not an absolute physiological ground truth. Across both tasks, I-VT generally showed high agreement with the benchmark and the most favorable balance between fixation and saccade classification. Task structure influenced the optimal thresholds obtained by $K$-ratio minimization. The Waldo dataset required higher velocity and angular-velocity thresholds than the Random Pixel dataset, which reflects more variable and goal-directed oculomotor dynamics in visual search. At the same time, both datasets were dominated by fixations, with saccades forming only a small fraction of samples. This class imbalance means that accuracy alone is not a sufficient measure of performance. This is evident in noise analysis, where adaptive I-DT maintains high overall accuracy largely by preserving fixation performance, while saccade detection (particularly recall) degrades. The class-wise metrics in the Results show that a classifier can achieve relatively high accuracy while failing to identify many saccades. This supports the use of F1-scores, confusion matrices, and explicit event proportions, in line with earlier discussions of evaluation practice in eye movement classification [zemblys].
@@
The $K$-ratio criterion provided a simple, task-specific method to tune thresholds without explicit modelling of the underlying velocity distributions. By minimizing the discrepancy between empirical and independent transition probabilities, it favored segmentation in which fixations and saccades formed temporally coherent states. This approach is complementary to more elaborate state-space models, such as hidden Markov models, which require stronger assumptions about observation statistics [10.1145/355017.355028]. Here, the $K$-ratio produced consistent participant-specific thresholds and offered a transparent way to adapt classical I-VT, I-AVT, and I-DT schemes to raw eye-tracking data. The analysis of individual differences supports the adaptive interpretation of the method: the $K$-ratio minimum is stable enough to estimate at the participant level, yet variable enough across participants that individualized threshold selection is justified. Notably, the repeatability of the minimum $K$-ratio itself was consistently higher than the repeatability of the threshold location. This indicates that the optimization criterion is robust even when the exact threshold achieving the minimum shifts somewhat across trials.
@@
The noise analysis confirmed that all three algorithms are sensitive to measurement noise. Even modest Gaussian perturbations led to a marked reduction in agreement with the EyeLink labels and to substantial shifts in event composition. In all cases, re-optimizing thresholds by $K$-ratio mitigated, but did not entirely prevent, the loss of performance. These patterns are consistent with the intended effect of transition minimization, which stabilizes segmentation under noise by prioritizing temporal coherence, but can reduce sensitivity to short saccades. These findings agree with previous work emphasizing the importance of careful preprocessing and denoising in eye-tracking analysis [niehorster2020noise, niehorster2017denoising, LARSSON2015145]. From an applications perspective, the results suggest that I-VT can serve a reasonable default when high-quality data are available, and the goal is to obtain a reliable first-order segmentation of fixations and saccades. I-AVT, despite its directional component, offered limited benefit in this context because of its strong under-detection of saccades. I-DT may still be useful in settings where long, stable fixations are of primary interest, such as some reading and static-scene paradigms [68d3bfb814cc4a61aae0131dd135c9ec], but it is less suitable for dynamic visual search. In all cases, practitioners should be aware that fixed thresholds calibrated on one task or recording setup may not transfer to others, and that auxiliary events such as microsaccades, smooth pursuits, and glissades remain largely unmodeled [martinezs].
@@
There are several natural directions for future work. One is to combine $K$-ratio tuning with more advanced denoising and interpolation schemes, which could extend the useful operating range of classical algorithms in noisy or mobile settings. Another is to embed $K$-ratio or related transition statistics into learning-based classifiers, for example, as regularization terms or diagnostic tools for model calibration [FRANCESCHIELLO2022106929, zemblys]. A third is to pursue more systematic study of inter-individual differences, with thresholds optimized at the individual level, to clarify how much of the observed variability is due to genuine oculomotor differences and how much is an artifact of algorithm choice [HESSELS2019100710]. Because both our methods and the EyeLink parser are rule-based detectors, the reported accuracies quantify agreement with a standard commercial parser, which validates against expert manual coding [Andersson2017OneAlgorithm] or synthetic ground-truth data remains an important direction for future work.
@@
Competing interests
The authors declare no competing interests.
@@ Acknowledgements
Appendix A Assumption checks for the K-ratio framework
To examine whether the stationarity assumption used to express the $K$-ratio in empirical form (Eq. (7)) is reasonable for our tasks, we performed two diagnostics on the reference fixation: saccade label sequence. First, we divided each recording of the Random Pixel task (3 trials, each of 60 seconds) into 12 equal-duration, non-overlapping temporal blocks, and computed within each block the saccade fraction $n_{S}$, the one-step transition probabilities $p(F\to S)$ and $p(S\to F)$, and the corresponding blockwise $K$-ratio. This provides a direct check for slow within-task drift in event composition and switching. In the Random Pixel example shown in Fig. 5, $n_{S}$ varies modestly across blocks (0.0730.0730.1660.166). At the same time, transition probabilities remain close ($p(F\to S)\in[0.00182,0.00294]$ and $p(S\to F)\in[0.00182,0.00294]$ per step). The $K$-ratio stays in a narrow range ($K\in[0.021,0.034]$, mean $0.026\pm 0.003$), indicating that the transition-minimization criterion is stable under mild nonstationarity. Second, we evaluated short-lag Markov adequacy by comparing empirical $k$-step transition matrices to the first-order Markov prediction $T^{k}$. The mean absolute deviation $\|T_{\mathrm{emp}}(k)-T^{k}\|$ is near zero at short lags (e.g., $3.35\times 10^{-4}$ at $k=2$, $9.98\times 10^{-4}$ at $k=3$, and $3.21\times 10^{-3}$ at $k=5$), supporting a first-order proxy at the timescale most relevant for suppressing spurious switching and selecting thresholds via $K$-ratio minimization. These checks do not assert that gaze dynamics are globally stationary; rather, they show that a first-order Markov approximation is an adequate local model for the purpose of threshold tuning in our datasets.
Figure 5: Stationarity and Markov adequacy diagnostics. (a) Blockwise $n_{S}$, $p(F\to S)$, $p(S\to F)$. (b) Blockwise (K)-ratio stability. (c) Deviation $|T_{\mathrm{emp}}(k)-T^{k}|$ as a function of lag $k$, showing very small error at short lags.
In addition to the visual blockwise diagnostics, we performed a regression-based trend check across blocks. For $n_{S}$, the linear trend was not significant ($p=0.556$), and the quadratic component was marginal ($p=0.060$), while for the $K$-ratio, neither a linear trend ($p=0.278$) nor a quadratic component ($p=0.941$) was significant. This supports the interpretation that any slow drift is modest and does not invalidate a first-order Markov proxy for local threshold tuning.
Appendix B Stability of participant-specific threshold, and $K$-ratio across trials
Figure 6: Distribution of participant-specific optimal thresholds obtained by minimizing the $K$-ratio, shown separately for I-VT, I-AVT, and I-DT. Gray open circles represent trial-level optimal thresholds, blue filled circles indicate participant-wise mean thresholds across trials, and red horizontal lines mark the grand mean across all observations for each method.
To examine whether the $K$-ratio minimum reflects a stable participant-specific property rather than only a pooled task-level optimum, we repeated the threshold optimization separately for each participant and each trial. For the Random Pixel task, optimization was performed for the three 60 s trials. For the Waldo task, optimization was performed for the nine 45 s trials. This yielded, for each participant and each trial, an optimal threshold for I-VT, I-AVT, and I-DT together with the corresponding minimum $K$-ratio value.
Figure 7: Trial-wise participant-specific thresholds obtained by minimizing the $K$-ratio. Top row: Random Pixel. Bottom row: Waldo. Panels show I-VT, I-AVT, and I-DT separately. Thin colored lines correspond to individual participants. The thick dark line shows the mean threshold across participants at each trial, and the shaded region indicates the standard error of the mean.
The resulting trial-wise thresholds are shown in Fig. 6 and Fig. 7. For each measure $y_{ij}$, where $i$ indexes participant and $j$ trial, repeatability across trials was quantified using the one-way random-effects intraclass correlation coefficient, ICC(1,1)(1,1). Under the variance-components model
$y_{ij}=\mu+a_{i}+\varepsilon_{ij},$ (11)
with between-participant variance $\sigma_{p}^{2}$ and within-participant variance $\sigma_{t}^{2}$, the ICC is
$\mathrm{ICC}=\frac{\sigma_{p}^{2}}{\sigma_{p}^{2}+\sigma_{t}^{2}}.$ (12)
In practice, ICC(1,1)(1,1) was estimated from the ANOVA mean squares as
$\mathrm{ICC}(1,1)=\frac{MS_{\mathrm{between}}-MS_{\mathrm{within}}}{MS_{\mathrm{between}}+(k-1)MS_{\mathrm{within}}},$ (13)
where $MS_{\mathrm{between}}$ and $MS_{\mathrm{within}}$ are the between- and within-participant mean squares, respectively, and
$k=\frac{n}{\sum_{i=1}^{n}1/k_{i}}$ (14)
is the harmonic mean number of valid trials per participant for the unbalanced design. Statistical significance was assessed with the standard ANOVA ratio
$F=\frac{MS_{\mathrm{between}}}{MS_{\mathrm{within}}},$ (15)
with degrees of freedom $df_{1}=n-1$ and $df_{2}=N-n$, where $n$ is the number of participants and $N=\sum_{i}k_{i}$ the total number of observations.
The ICC results are summarized in Table 2. In the Random Pixel task, the participant-specific optimal thresholds were moderately repeatable across trials for I-VT and I-AVT, with ICC values of 0.688 and 0.676, respectively. In contrast, the I-DT spatial threshold showed substantially weaker repeatability, with ICC = 0.257. The minimum $K$-ratio values at the participant-specific optima were more repeatable than the thresholds themselves, with ICC values of 0.816 for I-VT, 0.808 for I-AVT, and 0.779 for I-DT.
In the Waldo task, repeatability was generally stronger. The I-VT threshold yielded ICC = 0.728, the I-AVT threshold ICC = 0.544, and the I-DT spatial threshold ICC = 0.808. The minimum $K$-ratio values again showed high repeatability, with ICC values of 0.853 for I-VT, 0.845 for I-AVT, and 0.859 for I-DT. All Waldo ICC estimates were highly significant. In the Random Pixel task, all intraclass correlations were significant except for the case of the I-DT threshold, which was only marginal ($p=0.050$).
Figure 8: Participant-specific optimal thresholds estimated separately for each trial by minimizing the $K$-ratio. Colored dots represent trial-level values and black open circles represent participant means.
These findings support two related conclusions. First, the threshold that minimizes the $K$-ratio is not merely a pooled artifact. It shows measurable within-participant consistency across trials, indicating that the optimization is capturing stable structure in the eye-movement sequence. Second, the thresholds are not identical across participants. This between-participant variation is visible in Fig. 7 and is reflected statistically by high ICC values.
Table 2: Repeatability across trials of participant-specific optimal thresholds and corresponding minimum $K$-ratio values, quantified using one-way random-effects intraclass correlation, ICC(1,1)(1,1). A higher ICC indicates greater within-participant consistency across trials.
Task Measure ICC(1,1)(1,1) $F$ $p$
RPixel I-VT threshold 0.688 7.613 $3.92\times 10^{-6}$
RPixel I-AVT threshold 0.676 7.257 $6.24\times 10^{-6}$
RPixel I-DT threshold 0.257 2.036 0.0502
RPixel $K$-ratio at I-VT optimum 0.816 14.296 $4.58\times 10^{-9}$
RPixel $K$-ratio at I-AVT optimum 0.808 13.621 $7.97\times 10^{-9}$
RPixel $K$-ratio at I-DT optimum 0.779 11.603 $1.64\times 10^{-8}$
Waldo I-VT threshold 0.728 21.695 $10^{-15}$
Waldo I-AVT threshold 0.544 10.216 $2.50\times 10^{-14}$
Waldo I-DT threshold 0.808 33.854 $10^{-15}$
Waldo $K$-ratio at I-VT optimum 0.853 45.729 $10^{-15}$
Waldo $K$-ratio at I-AVT optimum 0.845 43.072 $10^{-15}$
Waldo $K$-ratio at I-DT optimum 0.859 48.274 $10^{-15}$
Appendix C Noise robustness of fixed and adaptive thresholds
@@
In this supplementary section, we provide detailed results from the noise robustness analysis. Table 3 reports the task-specific thresholds obtained by $K$-ratio minimization under noise-free conditions. These values served as the fixed (non-adaptive) baseline thresholds in subsequent experiments.
@@
Table 3: $K$-ratio-optimized thresholds for the three algorithms (zero added noise).
@@
Table 4: Classification degradation with fixed thresholds under increasing noise for I-VT algorithm. Note: Agr. = agreement with noise level 0 ($\sigma=0$) (%); Fix/Sac = sample counts; $\Delta$% = change in fixation proportion from baseline.
@@
Table 5: I-VT: Original vs. Adaptive threshold performance across noise levels. Note: Acc = Accuracy; F1F = F1-score (Fixation); F1S = F1-score (Saccade).
@@
Table 6: I-AVT: Original vs. Adaptive threshold performance across noise levels. See Table 5.
@@
Table 7: I-DT: Original vs. Adaptive threshold performance across noise levels. See Table 5.
@@
Figure 9: F1 scores for fixation and saccade classification across noise levels. Top row: F1-score for fixation detection. Bottom row: F1-score for saccade detection. Solid lines indicate original thresholds; dashed lines indicate adaptive thresholds.
Figure 10: Precision and recall for fixation detection across noise levels. Top row: Precision. Bottom row: Recall. Adaptive thresholds maintain higher precision at the cost of reduced recall at extreme noise levels.
Table 4 shows how classification performance degrades when optimal noise-free thresholds are held fixed under increasing Gaussian spatial noise ($\sigma=0$5050 px) for the I-VT algorithm. Agreement with the noise-free classification drops below 90% at $\sigma=2$ px for Random Pixel, while Waldo maintains agreement above 90% until $\sigma=5$ px. At high noise levels ($\sigma\geq 30$ px), both tasks approach chance-level agreement ($\sim$5–7%), with nearly all samples misclassified as saccades.
Tables 5-7 compare accuracy and F1 scores between fixed thresholds (optimized at $\sigma=0$) and adaptive thresholds (re-optimized at each noise level) for all three algorithms across both tasks. I-DT with adaptive thresholds demonstrates the greatest robustness, maintaining accuracy above 0.81 even at $\sigma=50$ px, whereas I-AVT shows the steepest degradation, falling to 0.31 accuracy at extreme noise levels.
Figure 9 visualizes F1 scores for fixation and saccade detection across noise levels. Adaptive thresholds (dashed lines) preserve fixation F1 scores substantially better than fixed thresholds (solid lines), with I-DT maintaining fixation F1 >> 0.90 across all noise levels. However, saccade F1 scores decline for all algorithms under both threshold conditions, with adaptive I-DT showing particularly low saccade F1 (<< 0.11) at high noise. Figure 10 reveals the precision-recall trade-off underlying adaptive threshold optimization. Precision for fixation detection remains high across noise levels under adaptive thresholding, but recall decreases substantially at $\sigma>10$ px. This pattern indicates that adaptive thresholds achieve overall accuracy by conservatively classifying uncertain samples as fixations.